# On the Origin of the High Star-Formation Efficiency in Massive Galaxies at Cosmic Dawn

### v1

@@ 3.5 Star formation histories
At a high level, we describe the SFH using summary statistics in Table 1, including the total stellar mass $M_{*}$ in the galaxy at $z=9$, the average SFR over the last $50~{}{\rm Myr}$ at $z=9$ ${\rm SFR}_{50}$, the integrated SFE $\epsilon_{\rm int}$, the median local SFE ${\rm med}(\epsilon_{\rm ff})$, and the outflow efficiency $\eta$. The calculation of these summary statistics is described in Appendix D. Nearly all our simulations produce significant stellar populations ($M_{*}\gtrsim 10^{9}~{}M_{\odot}$) with high SFRs ($\dot{M}_{*}\gtrsim 100~{}M_{\odot}/{\rm yr}$) and high SFEs ($\epsilon_{\rm ff},\epsilon_{\rm int}\gtrsim 10\%$), supporting the notion that star formation is intrinsically efficient at Cosmic Dawn.
Throughout this paper, we use 3 different measurements of SFE. The local SFE $\epsilon_{\rm ff}$ is the fraction of gas within a single cell that is converted into stars within one local freefall time. The global SFE $\epsilon_{\rm glob}=\dot{M}_{*}/f_{\rm b}\dot{M}_{\rm acc}$ is the ratio of the SFR in the galaxy to the gas accretion rate. The integrated SFE $\epsilon_{\rm int}=M_{*}/f_{\rm b}M_{\rm halo}$ is the ratio of the stellar mass to the total gas mass accreted onto the galaxy over its lifetime. The distinction between these SFEs highlights that any measure of SFE is defined on a particular spatial and temporal scale. Measurements of the SFE on multiple scales are necessary ingredients of a holistic picture of star formation.
Figures 9 and 10 visually represent the integrated and local SFEs respectively. In Figure 9, we show the stellar mass as a function of time, which is proportional to the integrated SFE. In Figure 10, we show the distribution of local SFEs in star formation events.
@@
In yesFbk, both feedback mechanisms operate simultaneously. yesFbk has the lowest integrated and local SFE in the feedback series. The low integrated SFE is surprising because it suggests that photoionization is effective, but only in tandem with SN feedback. The low local SFE is also surprising, because it suggests that when both feedback mechanisms work together, they alter the local gas conditions, even though this effect is not seen for either mechanism in isolation.
@@
As photoionization feedback becomes more efficient, the outflow efficiency increases, reaching $\eta\simeq 10$ in highPhot. This suggests that photoionization feedback suppresses star formation by enhancing the ability of SN explosions to drive outflows. This interpretation is supported by the local gas properties, which show that photoionization feedback is associated with (i) an increase in the volume fraction of the WNM (Fig. 6) and (ii) more SN explosions in photoionized gas and HIM rather than CNM (Fig. 8).
When SNe occur in more diffuse gas, more of their energy makes it into the ISM rather than being absorbed locally and radiated away. Once in the ISM, this energy can suppress star formation by disrupting star-forming clouds and launching outflows.
These data do not directly prove our interpretation because they only establish correlation rather than causation. However, previous works have come to similar conclusions regarding the interplay of photoionization and SN feedback. For example, Rosdahl et al. (2015) find that photo-heating smooths the density field in their disc galaxies, enhancing the effectiveness of SN feedback (their Fig. 9).
@@ 3.6 Star formation variability
After major bursts, the SFR temporarily decreases, possibly because the bursts consume dense gas available to form stars. The movies associated with this paper reveal that the major bursts in the SFR coincide with significant merger events, although analyzing the role of mergers in detail is beyond the scope of this paper.
We quantify the variability of the SFR using the power spectral density (PSD) in the time domain. To compute the PSD, we follow a similar procedure to Iyer et al. (2020). We start with the SFR in time bins of width $\Delta t=0.1~{}{\rm Myr}$, which gives a reasonable balance between shot noise and time resolution. We work with the logarithm of the SFR so that fluctuations in the SFR are measured relative to the mean rather than on an absolute scale. If we worked with the raw SFR, then the variability calculated at early times would be smaller than the variability calculated at late times just because the mean SFR is lower.
@@
The transition between these two regimes occurs on a timescale $\sim 1~{}{\rm Myr}$, which can be interpreted as our star formation timescale i.e. the free-fall time in a star forming cloud. On shorter timescales, it is impossible for one star formation event to be correlated with another. A freefall time $\sim 1~{}{\rm Myr}$ is associated with a density $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$, which we can use as an estimate for the typical density of star formation.
On longer timescales, the PSD becomes flat again. The transition between these two regimes occurs on a timescale $\sim 10~{}{\rm Myr}$, which can be interpreted as the lifetime of star forming clouds in our simulation (Tacchella et al., 2020). Due to the 550 Myr duration of our simulations, we cannot compute the PSDs on longer timescales using Welch’s method. However, one might expect the PSDs to regain a positive slope at even longer timescales due to correlations caused by galaxy mergers, outflow recycling, and other processes affecting the gas reservoir (Tacchella et al., 2020).
@@ 3.7 Variable turbulence forcing parameter
In varTurb, we determine the turbulence forcing parameter from the local velocity field using Equation 15. varTurb behaves almost like a simulation with constant forcing parameter in between the solenoidal and compressive forcing limits. In Figure 7, the density, temperature, and turbulent velocity dispersion of star-forming cells for varTurb is somewhere in between the distributions for solTurb and compTurb.
@@ 4.2 Metal enrichment
This model assumes that metals produced in SN explosions are rapidly mixed throughout the ISM rather than remaining localized to the SN site. The turbulent environment at Cosmic Dawn facilitates that mixing. Our simulated galaxies have typical velocity dispersions $\sigma_{\rm turb}\simeq 40~{}{\rm km/s}$ and radii $R_{\rm gal}\simeq 1~{}{\rm kpc}$, so the metal mixing time is approximately $\tau_{\rm mix}\sim R_{\rm gal}/\sigma_{\rm turb,1D}\simeq 30~{}{\rm Myr}$, far shorter than the lifetime of the galaxy.
@@ 4.3 Condition for self-regulation
where $f_{\rm sf}=M_{\rm gas}/M_{\rm sf}$ is the star-forming mass fraction and $\mu\dot{M}_{*}$ is the rate at which gas mass is removed from the star-forming state by stellar feedback. $\mu$ characterizes the strength of preventative feedback.
@@
We can combine Equation 48 and our simulations to estimate the depletion time in MDGs. In our simulations, the typical local SFE is $\epsilon_{\rm ff}\sim 10\%$ and the typical density of star-forming regions is $n_{\rm H}=3\times 10^{3}~{}{\rm cm^{-3}}$, corresponding to a freefall time $\tau_{\rm ff}\sim 1~{}{\rm Myr}$. Feedback is weak, so we can approximate $\mu\approx 0$. $f_{\rm dyn}$ and $\tau_{\rm cool}/\tau_{\rm ff}$ are both factors of a few. These values imply that the depletion time is on the order of 10s of ${\rm Myr}$.
@@
For a halo of mass $M_{\rm halo}=10^{11}~{}{\rm M_{\odot}}$ at redshift $z=9$, we find $\tau_{\rm acc}\simeq 154~{}{\rm Myr}$. This is longer than the depletion time, so the assumption of steady state in Section 4.1 and the resulting self-regulation behavior are valid.
@@ 4.4 Turbulence
Making the turbulence more compressive also has a broadening effect, because $\mathcal{M}_{\rm turb}^{2}$ and $b_{\rm turb}$ play the same role in Equation 12. This explains the trend of increasing local SFE as turbulence becomes more compressive. However, changing the TKE has an additional effect which does not occur when changing the turbulence forcing parameter.
@@
In Figure 12, we show how the local SFE in the MFF model varies as a function of temperature, density, and turbulent velocity dispersion in the solenoidal and compressive forcing limits at a fixed smoothing scale $\Delta x=10~{}{\rm pc}$. Physically, the smoothing scale can be understood as a mixing length. In numerical contexts, it can be understood as the size of a resolution element.
@@
Figure 12: Local SFE as a function of temperature and Hydrogen number density in the MFF model (Eq. 22). Each column shows a different turbulent velocity dispersion. The top row is the solenoidal forcing limit ($b_{\rm turb}=0.3$) and the bottom row is the compressive forcing limit ($b_{\rm turb}=1.0$). The resolution scale is set to its minimum value $\Delta x=10~{}{\rm pc}$ in our simulations, and we emphasize that the local SFE is a resolution-dependent quantity by definition. Iso-efficiency contours are shown in red for $\epsilon_{\rm ff}\in\{0.1\%,1\%,10\%,100\%\}$. When $\epsilon_{\rm ff}\sim 0\%$, the gas is Jeans stable. When $\epsilon_{\rm ff}=100\%$, the gas is fully Jeans unstable. In each panel, we overplot a blue contour outlining the region of phase space containing 90% of star formation events weighted by stellar mass in solTurb (top row) and compTurb (bottom row). The contours are smoothed by a $\sigma=10~{}{\rm px}$ Gaussian filter for ease of viewing. In general, the local SFE decreases with increasing turbulent velocity dispersion for the same gas conditions, except for gas conditions which are marginally Jeans stable in the compressive forcing limit.
@@
Due to self-regulation (Sec. 4.1), the decrease in local SFE due to the high turbulence of Cosmic Dawn only has a weak effect on the stellar mass which ultimately forms. For turbulent velocity dispersions even larger than the typical values $\sigma_{\rm turb}\sim 40~{}{\rm km/s}$ in our simulated galaxies, the local SFE may drop enough to take the galaxy out of the self-regulated regime. In this case, the integrated SFE would begin to decrease in proportion to the decrease in local SFE.
@@ 4.5 Other early feedback processes
Our early feedback model only includes thermal pressure from photoionized gas (Sec. 2.5). However, early feedback processes also include FUV radiation pressure, multiply-scattered IR radiation pressure, and stellar winds. All of these processes can be treated self-consistently by coupling the hydrodynamics to a radiative transport solver, for example using RAMSES-RT (Rosdahl et al., 2013). However, this extra physics comes at a computational cost. In this work, we prioritized modeling a large parameter space and understanding the physics in the simplified case of pure hydrodynamics.
Previous results from RAMSES-RT give clues about how our results might be affected by a more realistic early feedback model. Rosdahl et al. (2015) use RAMSES-RT to study the effect of radiative processes in a Milky-Way-like galaxy. They find that photoionization-heating is the dominant form of early feedback. In their simulations, photoionization feedback suppresses star formation by preventing the formation of star-forming clouds rather than by destroying them.
In principle, IR radiation pressure feedback should become more important in the high-density environments of Cosmic Dawn. IR photons can be absorbed and reemitted by dust multiple times when the IR optical depth is high. Even though typical UV dust opacities are $\sim 100$ times greater than IR dust opacities (Semenov et al., 2003), the multiply-scattered IR photons collectively exert more radiation pressure than the singly-scattered FUV photons.
Rosdahl et al. (2015) admit that they may underestimate the impact of IR radiation pressure feedback because the densest gas clouds are unresolved by their $\Delta x_{\rm min}=10~{}{\rm pc}$ resolution. However, Menon et al. (2022) suggest that the effect IR radiation pressure feedback is mitigated by the presence of turbulence, even at high surface densities. The authors explain that in an inhomogeneous medium, photons preferentially escape through low-density channels, increasing the escape fraction. This matter-radiation anti-correlation has been previously discussed in the context of massive star formation (Krumholz et al., 2009; Rosen et al., 2016) and lower densities regimes (Skinner & Ostriker, 2015; Tsang & Milosavljević, 2018).
Matter-radiation anti-correlation might be even more important in MDGs compared to Milky-Way-like galaxies. Our simulations indicate that MDGs host strong turbulence $\mathcal{M}_{\rm turb}\sim 30$. The turbulence is particularly strong in the dense regions which contribute most to the optical depth. Strong turbulence provides more low-density channels through which radiation can escape. In addition, radiation pressure feedback might self-regulate by driving a Rayleigh-Taylor instability which enhances turbulent motions further (Krumholz & Thompson, 2012). Another consideration at Cosmic Dawn is the low dust-to-gas ratio, which may reduce the effectiveness of radiation pressure feedback (Menon et al., 2024).
In Figure 13, we show how these effect conspire to reduce the IR optical depth in our simulations. In the left panel, we do not account for matter-radiation anti-correlation. In the other panels, we account for matter-radiation anti-correlation using an effective optical depth in the solenoidal and compressive forcing limits (Appendix G).
In this calculation, we determine the IR opacity using $\kappa_{\rm IR}=\tilde{\kappa}_{\rm IR}(Z/Z_{\odot})$, where $\tilde{\kappa}_{\rm IR}$ is the IR opacity at solar metallicity. We have assumed that the dust-to-gas ratio scales linearly with metallicity, which may overestimate the opacity at low metallicity (Feldmann, 2015; Choban et al., 2022). The opacity to IR photons at solar metallicity is roughly constant $\tilde{\kappa}_{\rm IR}\sim 5~{}{\rm g/cm^{2}}$ in the dust temperature range $100~{}{\rm K}\lesssim T_{\rm dust}\lesssim 1200~{}{\rm K}$, and takes on smaller values outside of that range (Semenov et al., 2003), so we adopt this constant value for the opacity and ignore the temperature dependence.
We find that the galaxy is only optically thick to IR photons when we do not account for matter-radiation anti-correlation. This suggests that the low metallicities and strong turbulence in MDGs can effectively suppress radiation pressure feedback from multiply-scattered IR photons.

### v2

@@ 3.5 Star formation histories
At a high level, we describe the SFH using summary statistics in Table 1, including the total stellar mass $M_{*}$ in the galaxy at $z=9$, the average SFR over the last $50~{}{\rm Myr}$ at $z=9$ ${\rm SFR}_{50}$, the integrated SFE $\epsilon_{\rm int}=M_{*}/f_{\rm b}M_{\rm halo}$, the median local SFE ${\rm med}(\epsilon_{\rm ff})$, and the outflow efficiency $\eta=\dot{M}_{\rm out}/{\rm SFR}_{\rm 50}$, where $\dot{M}_{\rm out}$ is the mass outflow rate. The calculation of these summary statistics is described in Appendix D. Nearly all our simulations produce significant stellar populations ($M_{*}\gtrsim 10^{9}~{}M_{\odot}$) with high SFRs ($\dot{M}_{*}\gtrsim 100~{}M_{\odot}/{\rm yr}$) and high SFEs ($\epsilon_{\rm ff},\epsilon_{\rm int}\gtrsim 10\%$), supporting the notion that star formation is intrinsically efficient at Cosmic Dawn.
In Figure 9, we plot the time evolution of the gas mass, stellar mass, and virial mass for the fiducial model. The details of the calculation are described in Appendix D. We also plot the gas fraction $f_{\rm gas}=M_{\rm gas}/(M_{\rm gas}+M_{*})$. By design, the virial mass reaches $10^{11}~{}M_{\odot}$ at $z=9$. Initially, the galaxy is dominated by the gas supplied by accretion. As the galaxy grows, it begins to form stars efficiently. At $z\simeq 10$, gas accretion is balanced by star formation, causing the gas mass to plataeu and the gas fraction to drop to $40\%$ by $z=9$. This process will likely continue until the gas fraction reaches values $\lesssim 10\%$ similar to massive galaxies in the local Universe (Wiklind et al., 2019).
Figure 9: Gas mass (blue), stellar mass (orange), virial mass (green), and gas fraction (top panel) as a function of time for the fiducial model. These quantities are computed at 43 discrete data dumps. By design, the virial mass reaches $10^{11}~{}M_{\odot}$ at $z=9$. At $z\simeq 10$, gas accretion is balanced by star formation, causing the gas mass to plataeu and the gas fraction to drop to $40\%$ by $z=9$.
Throughout this paper, we use three different measurements of SFE. The local SFE $\epsilon_{\rm ff}$ is the fraction of gas within a single cell that is converted into stars within one local freefall time. The global SFE $\epsilon_{\rm glob}=\dot{M}_{*}/f_{\rm b}\dot{M}_{\rm acc}$ is the ratio of the SFR in the galaxy to the gas accretion rate. The integrated SFE $\epsilon_{\rm int}=M_{*}/f_{\rm b}M_{\rm halo}$ is the ratio of the stellar mass to the total gas mass accreted onto the galaxy over its lifetime. Any measure of SFE is defined on a particular spatial and temporal scale. Measurements of the SFE on multiple scales are necessary ingredients of a holistic picture of star formation. Even simulations with a fixed local SFE $\sim 1\%$ can exhibit high integrated SFEs (e.g. Ceverino et al., 2024).
Figures 10 and 11 visually represent the integrated and local SFEs respectively. In Figure 10, we show the stellar mass as a function of time, which is proportional to the integrated SFE. In Figure 11, we show the distribution of local SFEs in star formation events.
@@
In yesFbk, both feedback mechanisms operate simultaneously. yesFbk has the lowest integrated and local SFE in the feedback series. The low integrated SFE suggests that photoionization is effective, but only in tandem with SN feedback. The low local SFE suggests that when both feedback mechanisms work together, they alter the local gas conditions, even though this effect is not seen for either mechanism in isolation.
These results are consistent with previous simulations, which show that early feedback pressurizes the gas in star-forming regions, driving expansion which reduces the gas density (e.g. Ceverino et al., 2014; Rosdahl et al., 2015). When SNe occur in more diffuse gas, more of their energy makes it into the ISM rather than being absorbed locally and radiated away. Once in the ISM, this energy can suppress star formation by disrupting star-forming clouds and launching outflows. The detailed simulations of dense gas clouds in Walch & Naab (2015) show that $\sim 50\%$ more SN energy goes into the ISM when SN go off in previously photo-ionized regions.
@@
As photoionization feedback becomes more efficient, the outflow efficiency increases, reaching $\eta\simeq 10$ in highPhot. This supports the idea that photoionization feedback enhances the coupling between SN explosions and the ISM. It is also consistent with the local gas properties, which show that photoionization feedback is associated with an increase in the volume fraction of the WNM (Fig. 6) and more SN explosions in photoionized gas and HIM rather than CNM (Fig. 8).
@@ 3.6 Star formation variability
After major bursts, the SFR temporarily decreases, possibly because the bursts consume dense gas available to form stars. The movies associated with this paper reveal that the major bursts in the SFR coincide with significant merger events, which was also seen in the cosmological zoom-in simulations of Cosmic Dawn galaxies in the FirstLight project (Ceverino et al., 2018). However, analyzing the role of mergers in detail is beyond the scope of this paper.
We quantify the variability of the SFR using the power spectral density (PSD) in the time domain. We compute the PSD following the procedure of Iyer et al. (2020). We start by computing the SFR using time bins of width $0.1~{}{\rm Myr}$, which gives a reasonable balance between shot noise and time resolution. The SFR spans a large dynamic range, so we work with the logarithm to better quantify the relative strengths of fluctuations in the SFR.
@@
Our techniques are similar to the recent variability analysis of Pallottini & Ferrara (2023) for the Serra simulations with a few differences. Pallottini & Ferrara (2023) fit the mean SFR trend to a polynomial in log space rather than an exponential, which is necessary because their simulations extend to lower redshifts where the mean SFR deviates from the exponential trend. They also use a Lomb (1976)-Scargle (1998) periodogram analysis rather than a PSD analysis. The former has several benefits including better isolation of individual frequency contributions and straight-forward quantification of peak significance (Dome et al., 2024). In future variability analyses, we will prefer this technique.
@@
The transition between these two regimes occurs on a timescale $\sim 1~{}{\rm Myr}$, which can be interpreted as the timescale for star formation i.e. the free-fall time in a star-forming cloud. On shorter timescales, it is impossible for one star formation event to be correlated with another. A freefall time $\sim 1~{}{\rm Myr}$ is associated with a density $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$, which is consistent with the typical density of star formation in Figure 7.
On longer timescales, the PSD becomes flat again. The transition between these two regimes occurs on a timescale $\sim 10~{}{\rm Myr}$, which can be interpreted as the lifetime of star forming clouds in our simulation (Tacchella et al., 2020). Due to the 550 Myr duration of our simulations, we cannot compute the PSDs on longer timescales using Welch’s method. However, one might expect the PSDs to regain a positive slope at even longer timescales due to correlations caused by galaxy mergers, outflow recycling, and other processes affecting the gas reservoir (Tacchella et al., 2020). For example, Ceverino et al. (2018) identify bursts in the SFR on timescales of 100s of Myr.
@@
For clarity of visualization, do not show our constant SFE simulations in Figure 12. In general, we find that the SFRs in the constant SFE simulations have less power in their PSDs than the corresponding MFF simulations, consistent with previous work (Martin-Alvarez et al., 2023).
@@ 3.7 Variable turbulence forcing parameter
In varTurb, we determine the turbulence forcing parameter from the local velocity field using Equation 15. varTurb behaves almost like a simulation with constant forcing parameter in between the solenoidal and compressive forcing limits. In Figure 7, the density, temperature, and turbulent velocity dispersion distributions of star-forming cells for varTurb are somewhere in-between the distributions for solTurb and compTurb.
@@ 4.2 Metal enrichment
This model assumes that metals produced in SN explosions are rapidly mixed throughout the ISM rather than remaining localized to the SN site. The turbulent environment at Cosmic Dawn facilitates that mixing. Our simulated galaxies have typical velocity dispersions $\simeq 50~{}{\rm km/s}$ and radii $\simeq 700~{}{\rm pc}$, so the metal mixing time is approximately $\tau_{\rm mix}\sim R_{\rm gal}/\sigma_{\rm turb,1D}\simeq 14~{}{\rm Myr}$, far shorter than the lifetime of the galaxy.
@@ 4.3 Condition for self-regulation
where $f_{\rm sf}=M_{\rm sf}/M_{\rm gas}$ is the star-forming mass fraction and $\mu\dot{M}_{*}$ is the rate at which gas mass is removed from the star-forming state by stellar feedback. $\mu$ characterizes the strength of preventative feedback.
@@
We can combine Equation 50 and our simulations to estimate the depletion time in MDGs. In our simulations, the typical local SFE is $\epsilon_{\rm ff}\sim 10\%$ and the typical density of star-forming regions is $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$, corresponding to a freefall time $\tau_{\rm ff}\sim 1~{}{\rm Myr}$. Feedback is weak, so we can approximate $\mu\approx 0$. $f_{\rm dyn}$ and $\tau_{\rm cool}/\tau_{\rm ff}$ are both factors of a few. These values imply that the depletion time is on the order of 10s of ${\rm Myr}$.
@@
For a halo of mass $10^{11}~{}{\rm M_{\odot}}$ at redshift $z=9$, we find $\tau_{\rm acc}\simeq 154~{}{\rm Myr}$. This is longer than the depletion time, so the assumption of steady state in Section 4.1 and the resulting self-regulation behavior are valid.
@@ 4.4 Turbulence
Making the turbulence more compressive also has a broadening effect, because $\mathcal{M}_{\rm turb}$ and $b_{\rm turb}$ play the same role in Equation 12. This explains the trend of increasing local SFE as turbulence becomes more compressive. However, changing the TKE has an additional effect which does not occur when changing the turbulence forcing parameter.
@@
In Figure 13, we show how the local SFE in the MFF model varies as a function of temperature, density, and turbulent velocity dispersion $\sigma_{\rm turb}=c_{\rm s}\mathcal{M}_{\rm turb}$ in the solenoidal and compressive forcing limits at a fixed smoothing scale $10~{}{\rm pc}$. Physically, the smoothing scale can be understood as a mixing length. In numerical contexts, it can be understood as the size of a resolution element.
@@
Figure 13: Local SFE as a function of temperature and Hydrogen number density in the MFF model (Eq. 22). Each column shows a different turbulent velocity dispersion. The top row is the solenoidal forcing limit ($b_{\rm turb}=0.3$) and the bottom row is the compressive forcing limit ($b_{\rm turb}=1.0$). The resolution scale is set to its minimum value $\Delta x=10~{}{\rm pc}$ in our simulations, and we emphasize that the local SFE is a resolution-dependent quantity by definition. Iso-efficiency contours are shown in red for $\epsilon_{\rm ff}\in\{0.1\%,1\%,10\%,100\%\}$. When $\epsilon_{\rm ff}\sim 0\%$, the gas is Jeans stable. When $\epsilon_{\rm ff}=100\%$, the gas is fully Jeans unstable. In each panel, we overplot a blue contour outlining the region of phase space containing 90% of star formation events weighted by stellar mass in solTurb (top row) and compTurb (bottom row). The contours are smoothed by a $10~{}{\rm px}$ Gaussian filter for ease of viewing. In general, the local SFE decreases with increasing turbulent velocity dispersion for the same gas conditions, except for gas conditions which are marginally Jeans stable in the compressive forcing limit.
@@
Due to self-regulation (Sec. 4.1), the decrease in local SFE due to the high turbulence of Cosmic Dawn only has a weak effect on the stellar mass which ultimately forms. For turbulent velocity dispersions even larger than the typical values $\sim 50~{}{\rm km/s}$ in our simulated galaxies, the local SFE may drop enough to take the galaxy out of the self-regulated regime. In this case, the integrated SFE would begin to decrease in proportion to the decrease in local SFE.
@@ 4.5 Other early feedback processes
Our early feedback model only includes thermal pressure from photoionized gas (Sec. 2.5). However, there are other early feedback processes we do not model, including radiation pressure from FUV and multiply-scattered IR photons and stellar winds. In this section, we discuss the effect of these mechanisms on star formation in MDGs.
4.5.1 Radiation pressure feedback
Massive stars produce a significant flux of UV photons with energies below the Lyman limit i.e. FUV. In dense environments like giant molecular clouds, these photons are readily absorbed by dust grains, imparting their momentum onto the surrounding gas (Krumholz & Matzner, 2009; Fall et al., 2010; Murray et al., 2010; Sales et al., 2014). This radiation pressure is augmented by multiply-scattered Ly-$\alpha$ (Kimm et al., 2018) and reprocessed IR (Murray et al., 2010; Skinner & Ostriker, 2015) photons. The former is not as well-studied, so we will focus our discussion on FUV and IR radiaton pressure feedback.
FUV radiation pressure feedback is diminished in high surface densities environments because the radiation pressure force cannot overcome the gravity of the cloud (Fall et al., 2010; Murray et al., 2010). In addition, radiation pressure forces partially cancel out when FUV is absorbed on scales smaller than the typical separation between radiation sources (Menon et al., 2023). On the other hand, IR radiation feedback is enhanced in high surface density environments because each IR photon undergoes many scatterings (Thompson et al., 2005; Krumholz & Matzner, 2009; Murray et al., 2010). When $\Sigma_{\rm b}\gtrsim 3000~{}M_{\odot}/{\rm pc}^{2}$, IR radiation pressure feedback dominates over FUV radiation pressure feedback (Menon et al., 2023), even though dust opacities are $\sim 100$ times greater in the FUV than in the IR (Semenov et al., 2003). Therefore, we expect IR radiation feedback to dominate over FUV radiation feedback in MDGs.
There are several additional considerations which may affect the strength of IR radiation feedback in MDGs. First, IR wavelengths are much larger than interstellar grain radii, so the dust opacity varies as $\kappa\propto\nu^{2}$ following the electric dipole limit (Draine, 2011). This causes the mean photon frequency to decrease as stellar radiation diffuses through the dust, weakening the dust-radiation coupling (Reissl et al., 2018; Krumholz et al., 2019). Second, the low metallicity at Cosmic Dawn leads to a low dust-to-gas ratio, which reduces the effective IR opacity (Menon et al., 2024).
Third, in a turbulent medium, radiation preferentially escapes through low column density channels (Skinner & Ostriker, 2015; Tsang & Milosavljević, 2018; Menon et al., 2022). Radiation pressure feedback may even self-regulate by driving turbulence (Krumholz & Thompson, 2012). This argument only applies to the IR photons, which are produced diffusely by dust and thus sample the turbulent density structure. The FUV photons are produced by the massive stars themselves which are preferentially surrounded by high-density regions. Matter-radiation anti-correlation is especially important in the star-forming regions of MDGs, where the turbulent velocity dispersions are high $\sim 50~{}{\rm km/s}$ (Fig. 7).
In Figure 14, we estimate the face-on IR dust optical depth in our fiducial simulation given by integrating the contributions from each cell along the projection direction. In the left panel, we compute the contribution of each cell as $\rho\kappa_{\rm IR}$, where $\kappa_{\rm IR}=\widetilde{\kappa}_{\rm IR}(Z/Z_{\odot})$ is the IR dust opacity at solar metallicity assuming that the dust-to-gas ratio scales linearly with metallicity. This may overestimate the opacity at low metallicity (Feldmann, 2015; Choban et al., 2022).
$\widetilde{\kappa}_{\rm IR}$ is roughly constant $\sim 5~{}{\rm g/cm^{2}}$ in the dust temperature range $100~{}{\rm K}\lesssim T_{\rm dust}\lesssim 1200~{}{\rm K}$, and takes on smaller values outside of that range (Semenov et al., 2003). We adopt this constant value and ignore the temperature dependence. In the other panels, we compute the contribution of each cell using an effective optical depth accounting for the matter-radiation anti-correlation (App. G) for solenoidal and compressive turbulence forcing models.
We find that despite the low metallicity, the star-forming regions of the galaxy are optically thick to IR. However, accounting for the matter-radiation anti-correlation significantly reduces the optical depth. Turbulent Mach number is positively correlated with density in our simulations (Fig. 19), so the regions that contribute most to the optical depth are most strongly affected by matter-radiation anti-correlation. This estimate suggests that IR radiation pressure feedback may still be important in MDGs, but matter-radiation anti-correlation needs to be accounted for. This should be investigated by future work.